Check whether each of the following subsets of R 3 is linearly independent. i) {(1,2,3),(−1,1,2),(2,1,1)}. ii) {(3,1,2),(−1,−1,−3),(−4,−3,0)


1
Expert's answer
2021-06-28T03:50:47-0400

Let us check whether each of the following subsets of "\R^3" is linearly independent.


i) "\{(1,2,3),(−1,1,2),(2,1,1)\}"


Since "\begin{vmatrix}1 & 2 & 3\\ −1 &1 &2 \\ 2 & 1 & 1\end{vmatrix}=1+8-3-6-2+2=0", we conclude that the subset "\{(1,2,3),(−1,1,2),(2,1,1)\}" is linearly dependent.



ii) "\{(3,1,2),(−1,−1,−3),(−4,−3,0)\}"

Taking into account that "\begin{vmatrix}3 &1&2\\−1&−1&−3\\−4&−3&0\end{vmatrix}=0+6+12-8-0-27=-17\ne 0," we conclude that the subset "\{(3,1,2),(−1,−1,−3),(−4,−3,0)\}" is linearly independent.